OpenStudy (anonymous):

Line AB contains points A (-2, 6) and B (4, 5). The slope of line AB is (5 points) Select one: a. -6 b.-1/6 c.1/6 d.6

OpenStudy (anonymous):

@ganeshie8

ganeshie8 (ganeshie8):

use slope formula

OpenStudy (anonymous):

its either a or b

OpenStudy (anonymous):

@eliassaab

OpenStudy (anonymous):

@ganeshie8

ganeshie8 (ganeshie8):

you're right. show me ur work, il give u answer i promise :)

OpenStudy (anonymous):

when i have these types of problems and i have to use the slope formula and i see a negative on one of the numbers i always cancel out the positive answers

ganeshie8 (ganeshie8):

thats indeed a clever trick :) you have the formula, so why not use the formula... take the royal path?

OpenStudy (anonymous):

i know -6 though i think

ganeshie8 (ganeshie8):

A (-2, 6) and B (4, 5) ^ ^ \(y_2 - y_1 = 5-6 = -1\)

ganeshie8 (ganeshie8):

A (-2, 6) and B (4, 5) ^ ^ \(x_2 - x_1 = 4--2 = 6\)

OpenStudy (anonymous):

so b right

ganeshie8 (ganeshie8):

slope = \(\dfrac{y_2-y_1}{x_2-x1} = ? \)

ganeshie8 (ganeshie8):

Yup, b is right !

ganeshie8 (ganeshie8):

slope = \( \dfrac{y_2-y_1}{x_2-x1} = \dfrac{-1}{6}\)